OpenStudy (anonymous):

Need help. Rewrite the following expression with positive exponents. (39xy) ^-4/7(negative exponent 4.7)

OpenStudy (anonymous):

move the equation down as the divisor of 1

OpenStudy (anonymous):

don't understand

OpenStudy (anonymous):

\[(39x y)^{-\frac{47}{10}}=\frac{1}{39^4 \left(39^{7/10}\right) (x y)^{47/10}} \]